OpenStudy (grandmoffchris):

2. Solve each given equation and show your work. Tell whether it has one solution, an infinite number of solutions, or no solutions, and identify each equation as an identity, a contradiction, or neither. A 6x + 4x - 6 = 24 + 9x B 25 - 4x = 15 - 3x + 10 - x C 4x + 8 = 2x + 7 + 2x - 20

OpenStudy (jackthegreatest):

@GrandMoffChris simplify each equation and tell me what you get

OpenStudy (grandmoffchris):

@TheSmartOne

OpenStudy (grandmoffchris):

I dont know how to simplify

OpenStudy (jackthegreatest):

I see, A: 6x + 4x - 6 = 24 + 9x 2x - 6 = 24 + 9x -30 = 7x x =? how many solutions do you think for this

OpenStudy (jackthegreatest):

ok you think 3, what would they be

OpenStudy (grandmoffchris):

10x - 6 = 24 + 9x

OpenStudy (grandmoffchris):

+

OpenStudy (jackthegreatest):

alright lets start from the basic stuff a solution of the equation is what the variable you are trying to find equals in this case, x

OpenStudy (jackthegreatest):

the solution should not be a equation, it should be some number of some sort

OpenStudy (jackthegreatest):

For equation 1, there is only 1 solution, what would it be @GrandMoffChris

OpenStudy (grandmoffchris):

Hmm, wait x when its 7x its 7 x x right?

OpenStudy (grandmoffchris):

No I mean 7 times many times

OpenStudy (grandmoffchris):

it multiplies itself

OpenStudy (grandmoffchris):

7 x 7 x 7

OpenStudy (jackthegreatest):

look at my work above

OpenStudy (jackthegreatest):

7x = -30 what does x =?

OpenStudy (grandmoffchris):

37?

OpenStudy (jackthegreatest):